# Bootstrapping Estimators based on the Block Maxima Method

### v1

@@ 4 Bootstrapping block maxima estimators
Note that $\hat{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*}$ depends on $k$, which is suppressed from the notation. However, unlike $\mathbb{G}_{n,r}^{({\operatorname{cb}}),*}$, this process is inconsistent for both $\bar{\mathbb{G}}_{n,r}^{({\operatorname{sb}})}$ and $\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}$. Indeed, considering the case $d=q=1$ for simplicity, we prove in the appendix that
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*}h\mid\mathcal{X}_{n}\big{)},\mathcal{N}(\bm{0},\Sigma_{h}^{(k)})\Big{)}=o_{\mathbb{P}}(1),$ (4.2)
@@
where $\Sigma_{h}^{(k)}=\Sigma_{h}^{({\operatorname{sb}})}-\frac{2}{k}\int_{0}^{1}\xi\operatorname{Cov}(h(\bm{Z}_{1,\xi}),h(\bm{Z}_{2,\xi}))\,\mathrm{d}\xi.$ If $\operatorname{Var}(h(\bm{Z}))>0$, we have $\Sigma_{h}^{(k)}<\Sigma_{h}^{({\operatorname{sb}})}$ for any $k\in\mathbb{N}$ by Lemma 11.4, so in view of Theorem 3.2 the bootstrap process has a smaller asymptotic variance than needed for approximating $\bar{\mathbb{G}}_{n,r}^{({\operatorname{sb}})}$ or $\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}$. This circumstance is one of the main motivations for working with circular block maxima.
@@
In statistical applications, as discussed after Theorem 2.3, we do not want to bootstrap the estimation error of empirical means involving the (unobservable) $\bm{Z}_{r,i}$, but the estimation error of general statistics depending on the block maxima $\bm{M}_{r,i}$ itself. We follow the general setting of Remark 2.4, that is, $\hat{\bm{\theta}}_{n}={\bm{\varphi}}_{n}(\bm{M}_{r,1},\dots,\bm{M}_{r,n})$ is some estimator of interest for a target parameter ${\bm{\theta}}_{r}\in\mathbb{R}^{p}$. If similar equations/linearizations as in (2.7) and (2.8) can be shown to hold for the (unobservable) circmax-bootstrap counterpart $\hat{\bm{\vartheta}}_{n}^{({\operatorname{cb}}),*}={\bm{\psi}}_{n}(\bm{Z}_{1,r}^{({\operatorname{cb}}),*},\dots,\bm{Z}_{n,r}^{({\operatorname{cb}}),*})$, that is, if
$\hat{\bm{\theta}}_{n}^{({\operatorname{cb}}),*}-\hat{\bm{\theta}}_{n}^{({\operatorname{cb}})}$ $=A(\bm{a}_{r},\bm{b}_{r})(\hat{\bm{\vartheta}}_{n}^{({\operatorname{cb}}),*}-\hat{\bm{\vartheta}}_{n}^{({\operatorname{cb}})}),$ (4.3)
$\sqrt{\frac{n}{r}}(\hat{\bm{\vartheta}}_{n}^{({\operatorname{cb}}),*}-\hat{\bm{\vartheta}}_{n}^{({\operatorname{cb}})})$ $=\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}(h_{1},\dots,h_{q})^{\top}+o_{\mathbb{P}}(1)$ (4.4)
@@
then, loosely speaking, the distributional approximation $\mathbb{G}_{n,r}^{({\operatorname{cb}})}\bm{h}\approx_{d}(\mathbb{G}_{n,r}^{({\operatorname{cb}}),*}\bm{h}\mid\mathcal{X}_{n})$ from Theorem 4.1 carries over to the distributional approximation $\hat{\bm{\theta}}_{n}^{({\operatorname{cb}})}-{\bm{\theta}}_{r}\approx_{d}(\hat{\bm{\theta}}_{n}^{({\operatorname{cb}}),*}-\hat{\bm{\theta}}_{n}^{({\operatorname{cb}})}\mid\mathcal{X}_{n})$. Formally, we have the following result.
@@
Fix $k\in\mathbb{N}_{\geq 2}$ and assume that Conditions 2.1 and 2.2 are met. Suppose $\hat{\bm{\theta}}_{n}^{({\operatorname{mb}})}={\bm{\varphi}}_{n}(\bm{M}_{r,1}^{({\operatorname{mb}})},\dots,\bm{M}_{r,n}^{({\operatorname{mb}})})$ is some estimator of interest for a target parameter ${\bm{\theta}}_{r}\in\mathbb{R}^{p}$ such that (2.5), (2.6) and (2.8) (with $\bar{\mathbb{G}}_{n,r}=\bar{\mathbb{G}}_{n,r}^{({\operatorname{mb}})}$) is met for some function $\bm{h}=(h_{1},\dots,h_{q})^{\top}$ satisfying the conditions of Theorem 3.2. Moreover, assume that (4.3) and (4.4) is met. Then, if $\Sigma_{\bm{h}}^{({\operatorname{sb}})}$ is invertible, for ${\operatorname{mb}}\in\{{\operatorname{sb}},{\operatorname{cb}}\}$,
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\bm{\theta}}_{n}^{({\operatorname{cb}}),*}-\hat{\bm{\theta}}_{n}^{({\operatorname{cb}})}\mid\mathcal{X}_{n}\big{)},\mathcal{L}\big{(}\hat{\bm{\theta}}_{n}^{({\operatorname{mb}})}-{\bm{\theta}}_{r}\big{)}\Big{)}=o_{\mathbb{P}}(1).$
Proof.
By Equations (2.5), (2.6), (2.8) and (4.3), (4.4) we have
$\hat{\bm{\theta}}_{n}^{({\operatorname{mb}})}-{\bm{\theta}}_{r}$ $=A(\bm{a}_{r},\bm{b}_{r})\sqrt{\frac{r}{n}}\big{(}\bar{\mathbb{G}}_{n,r}^{({\operatorname{mb}})}\bm{h}+o_{\mathbb{P}}(1)\big{)}=:\sqrt{\frac{r}{n}}A(\bm{a}_{r},\bm{b}_{r})S_{n},$
$\hat{\bm{\theta}}_{n}^{({\operatorname{cb}}),*}-\hat{\bm{\theta}}_{n}^{({\operatorname{cb}})}$ $=A(\bm{a}_{r},\bm{b}_{r})\sqrt{\frac{r}{n}}\big{(}\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}},\ast)}\bm{h}+o_{\mathbb{P}}(1)\big{)}=:\sqrt{\frac{r}{n}}A(\bm{a}_{r},\bm{b}_{r})S_{n}^{*}.$
By Theorems 2.3 and 3.2 we have $S_{n}\rightsquigarrow\mathcal{N}_{q}(\bm{0},\Sigma_{\bm{h}}^{({\operatorname{sb}})})$. Theorem 4.1 implies $d_{K}(\mathcal{L}(S_{n}^{*}\mid\mathcal{X}_{n}),\mathcal{L}(S_{n}))=o_{\mathbb{P}}(1).$ The assertion then follows from Lemma 11.9. ∎
The results from Proposition 4.4 are sufficient for showing that basic bootstrap confidence intervals (Davison and Hinkley, , 1997) asymptotically hold their intended level. As a proof of concept, details on a specific example are given in the next section, see Corollary 5.6.
@@ 5 Application: bootstrapping the pseudo-maximum likelihood estimator for the Fréchet distribution
In this section, we provide details how the previous methods and results can be used for a specific important estimation problem. More precisely, we consider the univariate heavy tailed case, and restrict attention to the maximum likelihood estimator for parameters associated with a suitable version of the max-domain of attraction Condition 2.1.

### v2

@@ 4 Bootstrapping block maxima estimators
Note that $\hat{\mathbb{G}}_{n,r}^{[{\operatorname{s}}],*}$ depends on $k$, which is suppressed from the notation. However, unlike $\hat{\mathbb{G}}_{n,r}^{[{\operatorname{c}}],*}$, this process is inconsistent for both $\bar{\mathbb{G}}_{n,r}^{[{\operatorname{s}}]}$ and $\bar{\mathbb{G}}_{n,r}^{[{\operatorname{c}}]}$, for any $k\in\mathbb{N}$. An heuristic explanation is provided in Section D in the supplement, where we also formally prove that, considering the case $d=q=1$ for simplicity,
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\mathbb{G}}_{n,r}^{[{\operatorname{s}}],*}h\mid\mathcal{X}_{n}\big{)},\mathcal{N}(\bm{0},\Sigma_{h}^{(k)})\Big{)}=o_{\mathbb{P}}(1),$ (4.2)
@@
where $\Sigma_{h}^{(k)}=\Sigma_{h}^{[{\operatorname{s}}]}-\frac{2}{k}\int_{0}^{1}\xi\operatorname{Cov}(h(\bm{Z}_{1,\xi}),h(\bm{Z}_{2,\xi}))\,\mathrm{d}\xi.$ If $\operatorname{Var}(h(\bm{Z}))>0$, we have $\Sigma_{h}^{(k)}<\Sigma_{h}^{[{\operatorname{s}}]}$ for any $k\in\mathbb{N}$ by Lemma C.4 in the supplement, so in view of Theorem 3.2 the bootstrap process has a smaller asymptotic variance than needed for approximating $\bar{\mathbb{G}}_{n,r}^{[{\operatorname{s}}]}$ or $\bar{\mathbb{G}}_{n,r}^{[{\operatorname{c}}]}$. This circumstance is one of the main motivations for working with circular block maxima.
@@
In statistical applications, as discussed after Theorem 2.3, we do not want to bootstrap the estimation error of empirical means involving the (unobservable) $\bm{Z}_{r,i}$, but the estimation error of general statistics depending on the block maxima $\bm{M}_{r,i}$ itself. We follow the general setting of Remark 2.4, that is, $\hat{\bm{\theta}}_{n}={\bm{\varphi}}_{n}(\bm{M}_{r,1},\dots,\bm{M}_{r,n})$ is some estimator of interest for a target parameter ${\bm{\theta}}_{r}\in\mathbb{R}^{p}$. If similar equations/linearizations as in (2.7) and (2.8) can be shown to hold for the (unobservable) circmax-bootstrap counterpart $\hat{\bm{\vartheta}}_{n}^{[{\operatorname{c}}],*}={\bm{\psi}}_{n}(\bm{Z}_{1,r}^{[{\operatorname{c}}],*},\dots,\bm{Z}_{n,r}^{[{\operatorname{c}}],*})$, that is, if
$\hat{\bm{\theta}}_{n}^{[{\operatorname{c}}],*}-\hat{\bm{\theta}}_{n}^{[{\operatorname{c}}]}$ $=A(\bm{a}_{r},\bm{b}_{r})(\hat{\bm{\vartheta}}_{n}^{[{\operatorname{c}}],*}-\hat{\bm{\vartheta}}_{n}^{[{\operatorname{c}}]}),$ (4.3)
$\sqrt{\frac{n}{r}}(\hat{\bm{\vartheta}}_{n}^{[{\operatorname{c}}],*}-\hat{\bm{\vartheta}}_{n}^{[{\operatorname{c}}]})$ $=\hat{\mathbb{G}}_{n,r}^{[{\operatorname{c}}],*}(h_{1},\dots,h_{q})^{\top}+o_{\mathbb{P}}(1)$ (4.4)
@@
then, loosely speaking, the distributional approximation $\mathbb{G}_{n,r}^{[{\operatorname{c}}]}\bm{h}\approx_{d}(\mathbb{G}_{n,r}^{[{\operatorname{c}}],*}\bm{h}\mid\mathcal{X}_{n})$ from Theorem 4.1 carries over to the distributional approximation $\hat{\bm{\theta}}_{n}^{[{\operatorname{c}}]}-{\bm{\theta}}_{r}\approx_{d}(\hat{\bm{\theta}}_{n}^{[{\operatorname{c}}],*}-\hat{\bm{\theta}}_{n}^{[{\operatorname{c}}]}\mid\mathcal{X}_{n})$. Formally, we have the following result.
@@
Fix $k\in\mathbb{N}_{\geq 2}$ and assume that Conditions 2.1 and 2.2 are met. Suppose $\hat{\bm{\theta}}_{n}^{[{\operatorname{m}}]}={\bm{\varphi}}_{n}(\bm{M}_{r,1}^{[{\operatorname{m}}]},\dots,\bm{M}_{r,n}^{[{\operatorname{m}}]})$ is some estimator of interest for a target parameter ${\bm{\theta}}_{r}\in\mathbb{R}^{p}$ such that (2.5), (2.6) and (2.8) (with $\bar{\mathbb{G}}_{n,r}=\bar{\mathbb{G}}_{n,r}^{[{\operatorname{m}}]}$) is met for some function $\bm{h}=(h_{1},\dots,h_{q})^{\top}$ satisfying the conditions of Theorem 3.2. Moreover, assume that (4.3) and (4.4) are met. Then, if $\Sigma_{\bm{h}}^{[{\operatorname{s}}]}$ is invertible, for ${\operatorname{m}}\in\{{\operatorname{s}},{\operatorname{c}}\}$,
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\bm{\theta}}_{n}^{[{\operatorname{c}}],*}-\hat{\bm{\theta}}_{n}^{[{\operatorname{c}}]}\mid\mathcal{X}_{n}\big{)},\mathcal{L}\big{(}\hat{\bm{\theta}}_{n}^{[{\operatorname{m}}]}-{\bm{\theta}}_{r}\big{)}\Big{)}=o_{\mathbb{P}}(1).$
The results from Proposition 4.4 are sufficient for showing that basic bootstrap confidence intervals (DavHin97) asymptotically hold their intended level. As a proof of concept, details on a specific example are given in the next section, see also Corollary A.6 in the supplement.
@@ 5 Application: bootstrapping the pseudo-maximum likelihood estimator for the Fréchet distribution
We provide details how the previous methods and results can be used for a specific important estimation problem, namely maximum likelihood estimation for parameters associated with a univariate heavy-tailed version of the MDA Condition 2.1. An extended version of this section, including formal mathematical statements and proofs, can be found in the supplement, Section A.